OpenStudy (anonymous):

the limit as x approaches negative infinity the equation is (x-2)/((((x^2)+1))^.5)) i have no idea how to get the limit on this equation.

myininaya (myininaya):

\[\lim_{x \rightarrow - \infty}\frac{x-2}{\sqrt{x^2+1}}*\frac{\frac{1}{\sqrt{x^2}}}{\frac{1}{\sqrt{x^2}}}\] since x<0, then sqrt{x^2}=|x|=-x. \[\lim_{x \rightarrow - \infty} \frac{-1-\frac{2}{\sqrt{x^2}}}{1+\sqrt{\frac{1}{x^2}}}\] \[\lim_{x \rightarrow -\infty}\frac{-1+\frac{2}{x}}{1+\frac{1}{-x}}\]

myininaya (myininaya):

=-1